### v1

# Grids Often Outperform Implicit Neural Representations

Abstract
Implicit Neural Representations (INRs) have recently shown impressive results, but their fundamental capacity, implicit biases, and scaling behavior remain poorly understood. We investigate the performance of diverse INRs across a suite of 2D and 3D real and synthetic signals with varying effective bandwidth, as well as both overfitting and generalization tasks including tomography, super-resolution, and denoising. By stratifying performance according to model size as well as signal type and bandwidth, our results shed light on how different INR and grid representations allocate their capacity. We find that, for most tasks and signals, a simple regularized grid with interpolation trains faster and to higher quality than any INR with the same number of parameters. We also find limited settings where INRs outperform grids—namely fitting signals with underlying lower-dimensional structure such as shape contours—to guide future use of INRs towards the most advantageous applications. Code and synthetic signals used in our analysis are available at https://github.com/voilalab/INR-benchmark.
1 Introduction
Signal representation is fundamental to sensing and learning, especially for spatial and visual applications including computer vision and computational imaging [1, 2, 3, 4]. Recently, implicit neural representations (INRs) have shown promising performance in a range of imaging and inverse problems, yielding high perceptual quality with small memory footprint. However, their representation capacity, scaling behavior, and implicit biases are not well understood, limiting the impact and confidence with which they can be deployed. Our work aims to shed light on these properties of different representations through principled experiments, to guide practitioners toward the most suitable strategy for their specific datasets and use cases as well as inform development of future signal representations.
Modern signal representations rely on four main strategies, each with unique advantages and limitations. Interpolated grid-based representations [5, 6, 7, 8] are fundamentally continuous, enjoy well-behaved gradients for fast optimization, and inherit classical representation guarantees rooted in sampling theory [9]. However, they suffer the curse of dimensionality: for fixed resolution, model size grows exponentially with dimension. In contrast, truly discrete representations such as point clouds and surface meshes [10, 11, 12, 13] can use fewer parameters to represent a sparse signal, but often lack useful gradients and require heuristic discrete optimization strategies that are sensitive to initialization. INRs model continuous signals using neural networks to map input coordinates to output signal values [1, 14, 15, 16, 17, 18]. They can provide plausible and stable reconstructions even with very limited model size, but suffer slow optimization and poorly understood resolution and scaling behavior. Hybrid approaches merge the strengths of both grids and neural networks [19, 20, 21, 22, 23, 24, 25, 26] by learning grid-based features alongside a lightweight neural network decoder, but questions remain regarding their representation capacity and implicit biases.
We compare these representation methods under a comprehensive suite of conditions, including both synthetic signals (e.g., bandlimited noise, fractal structures) and real-world data (e.g., natural images, CT scans). We consider tasks including overfitting images/volumes, denoising, super-resolution, and tomography. Our contributions are threefold:
  1. 1.
    We quantify the capacity of state-of-the-art INRs and hybrid models by evaluating their performance stratified by signal bandwidth, signal type, and model size. For our 2D Bandlimited signal class (see Figure 1), we observe that most models exhibit a power law relationship between model size and bandwidth.
  2. 2.
    We identify diverse scenarios, including overfitting 2D and 3D Bandlimited signals and solving inverse problems with both synthetic and real signals, in which simple uniform grids with interpolation offer pareto-optimal memory, speed, and reconstruction quality.
  3. 3.
    We identify specific scenarios where INRs and hybrid models can outperform grid-based representations, namely signals with underlying lower-dimensional structure, such as shape occupancy masks.
2 Methods
Spheres
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Sierpinski
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Band- limited
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
0.1 0.2 0.3 0.4 0.5 0.6 0.7 0.8 0.9
Figure 1: Synthetic Signals. Rows represent synthetic signal types, and columns represent effective bandlimits from 0.1 to 0.9. Signal detail and complexity increase with effective bandlimit.
Our experiments are designed to systematically assess diverse signal representations on both synthetic and real-world signals. We aim to address key gaps in understanding by evaluating each method’s ability to model signals with varying frequency content, generalize and solve inverse problems in 2D and 3D, and trade off parameter efficiency, representation power, and computation time. Specifically, we evaluate each model’s:
  1. 1.
    Expressivity: Ability to accurately overfit signals with varying feature scales within a fixed parameter budget.
  2. 2.
    Scalability: How performance changes with model size, particularly in terms of effective resolution and representation capacity.
  3. 3.
    Computation Time: Training and inference speeds, with a focus on how these scale with model size.
  4. 4.
    Inverse Problem Performance: Generalization performance for tasks such as super-resolution, denoising, and computed tomography (CT) reconstruction.
By systematically comparing representative grid-based, INR, hybrid, discrete models, we provide practical guidance for selecting appropriate representations for different application-specific needs.
Table 1: Summary of Synthetic and Real Signals
Dataset Type Dimensions Description
Synthetic Signals with “Bandlimits”
Bandlimited Signals 2D, 3D Random noise filtered radially in the Fourier domain with exponentially-spaced frequency cutoffs.
Spheres 2D, 3D Randomly arranged disks or balls that vary in size and number to represent features at different scales.
Sierpinski 2D Iterative fractal patterns with increasing structural complexity and fine detail.
Star Target 2D Radial triangular wedges, such that feature scale increases radially from the center.
Real Signals
DIV2K Images         [27] 2D High-resolution image dataset for (1) image overfitting, (2) 4×4$\times$4 × super-resolution along each axis, and (3) denoising with Gaussian noise (standard deviation $\epsilon\in\{0.05,0.1\}$).
Computed Tomography (CT) [28] 2D X-ray CT scan of a human chest, used to evaluate signal recovery in a classic underdetermined inverse problem.
3D Dragon Shape                  [29] 3D A solid 3D object with approximately $1\times 10^{6}$ voxels (before super-resolution). Models are evaluated for (1) overfitting and (2) super-resolution (doubling the total number of voxels).
3D Dragon Surface                 [29] 3D The surface of a 3D object with approximately $1\times 10^{6}$ voxels (before super-resolution). Models are evaluated for (1) overfitting and (2) super-resolution (doubling the total number of voxels).
2.1 Models
A broad overview of strategies for 2D and 3D signal representation in inverse problems is provided in Section 5.1. In our experiments, we compare eight representative approaches including pure INRs (Fourier Feature Networks [15], SIREN [14], WIRE [16]), hybrid methods (GA-Planes [26], Instant-NGP [22]), explicit representations (Gaussian Splatting [10] in 2D, via GSplat [30]), INRs with adaptive bandwidth (BACON [17]), and pure grids with interpolation (similar to [5]). Each model is tested across an exponential sweep of model sizes, with the number of trainable parameters ranging from $1\times 10^{4}$ to $3\times 10^{6}$ on signals with typical underlying dimension of $1\times 10^{6}$. All models are tuned by optimizing hyperparameters on our Star Target image (see row 4 in Figure 2). Further implementation details are provided in Section 5.5, Table 2, and Table 4 ; the code for both signal generation and model evaluation is provided at https://github.com/voilalab/INR-benchmark.
Fourier Feature Networks (FFN)
[15] embed the input coordinates $\mathbf{x}\in[-1,1)^{d}$ into $2m$ Fourier features, where $m=1000$ in our experiments. The Fourier feature mapping is defined as
$\gamma(\mathbf{x})=\begin{bmatrix}a_{1}\cos(2\pi\mathbf{\omega}_{1}^{T}\mathbf{x})&a_{1}\sin(2\pi\mathbf{\omega}_{1}^{T}\mathbf{x})\\ \vdots&\vdots\\ a_{m}\cos(2\pi\mathbf{\omega}_{m}^{T}\mathbf{x})&a_{m}\sin(2\pi\mathbf{\omega}_{m}^{T}\mathbf{x})\end{bmatrix},$
where the frequency vectors $\mathbf{\omega}_{i}$ are drawn from an isotropic multivariate Gaussian distribution with standard deviation $\sigma$ that tunes the bandwidth of the representation. In our implementation, these Fourier features are then decoded by a 2-hidden-layer ReLU MLP whose hidden dimension is scaled to vary model size while keeping the Fourier features unchanged.
SIREN
[14] uses a coordinate multilayer perceptron (MLP) with sinusoidal activation functions to represent smooth signals at varying scales. The elementwise activation function is:
$\psi(x;\omega)=\sin(\omega x),$
where the scalar hyperparameter $\omega$ tunes the bandwidth of the representation. While the original paper sets $\omega=30$ as the default, we use $\omega=90$ based on tuning on our Star Target image. Based on the original paper, we use a 3-hidden-layer SIREN and adjust the hidden dimension to control the model size.
WIRE
[16] uses a coordinate MLP with wavelet activation functions to capture multi-scale signal details that are localized in both space and frequency. The elementwise activation function is:
$\psi(x;\omega,s)=e^{j\omega x}e^{-|sx|^{2}},$
where the hyperparameters $\omega$ and $s$ control the frequency and spatial spread (width) of the wavelet. Based on the original paper, we use a 2-hidden-layer WIRE network and vary its hidden dimension to adjust the model size. We use $\omega=15$ and $s=10$ based on tuning on our Star Target image.
GA-Planes
[…]
$e_{c}=\psi(g_{c},\pi_{c}(\mathbf{q})),$
where $g_{c}$ is the feature grid for dimension tuple $c$, $\pi_{c}$ extracts the relevant coordinates from $\mathbf{q}$, and $\psi$ performs (bi/tri)linear interpolation. These interpolated features are then combined by multiplying combinations that yield the appropriate output dimension (line $\times$ line for 2D signals, and line $\times$ line $\times$ line and line $\times$ plane for 3D signals) and then these combined features are summed and decoded by a 2-layer ReLU MLP. The model size is scaled by varying the resolution and feature dimensions of the grids as well as the hidden dimension of the decoder MLP, following the general strategy recommended in the original paper to prioritize high resolution in lower-dimensional grids.
Instant-NGP
[22] uses a multi-resolution hash table encoding to represent 2D and 3D signals efficiently in both time and memory. For each resolution $l\in\{1,\dots,L\}$, the hash encoding is defined as:
$\mathbf{e}_{l}(\mathbf{x})=\psi\left(\mathbf{H}_{l}[\mathrm{hash}(\mathbf{p}_{l}(\mathbf{x}))]\right),$
where $\mathbf{p}_{l}(\mathbf{x})$ scales the input coordinate $\mathbf{x}$ to the resolution of the $l$-th grid, $\mathrm{hash}(\cdot)$ maps grid indices to feature table indices (allowing collisions), $\mathbf{H}_{l}$ is the hash table for resolution $l$, and $\psi$ performs (bi/tri)linear interpolation. Following the original paper implementation, the hash features from each resolution are concatenated and decoded by a 2-layer ReLU MLP with 64 neurons at each layer. We adjust model size by varying the length of the hash tables, while the MLP decoder remains fixed.
Gaussian Splatting (GSplat).
[10, 30] represents a scene with $N$ Gaussians $\mathcal{G}_{i}=(\boldsymbol{\mu}_{i},\mathbf{\Sigma}_{i},\mathbf{c}_{i},\alpha_{i})$. Each Gaussian is projected to screen space and rendered as a 2D density; the splats are then alpha-blended front-to-back in a fully differentiable rasterizer. Because this model does not provide a mechanism for direct 3D fitting, we restrict our evaluation of Gaussian Splatting to 2D signals, using the open-source gsplat implementation [30] to fit 2D Gaussians directly in image space.
Band-Limited Coordinate Networks (BACON)
[17] is a coordinate-based network that applies fixed sine filters at each layer. Each layer samples a frequency range $[-B_{i},B_{i}]$ and a phase shift, resulting in a cumulative bandwidth $\sum_{j=0}^{i}B_{j}$ after $i$ layers. Inputs are normalized to $[-0.5,0.5]^{d}$ and the signal is treated as periodic, allowing only a discrete set of frequencies to be represented.
Basic Interpolated Grid
directly stores parameter values at discrete lattice points in a grid, represented as $\mathbf{g}\in\mathbb{R}^{n_{1}\times n_{2}\times\cdots\times n_{d}}$, where $n_{i}$ represents the resolution of the grid along the $i$-th dimension. For a query point $\mathbf{q}\in\mathbb{R}^{d}$, the value is computed through interpolation between neighboring grid points. If the output is vector valued (e.g., RGB color channels) this is accommodated by adding an output dimension to the grid. Our 2D experiments use bicubic interpolation, which weights the 16 nearest grid values to produce a continuous and smooth (differentiable) function over 2D space. For 3D, we use trilinear interpolation which weights the nearest 8 grid values to produce a continuous but nonsmooth 3D function. Higher-order interpolation is also possible in 3D if a smooth function is desired, at higher computational cost. Further details on grid interpolation are provided in Section 5.2. We adjust model size by varying the grid resolution hyperparameters $n_{i}$, as the total model size scales with $\prod_{i=1}^{d}n_{i}$.
FFN SIREN WIRE GA-Planes Instant-NGP GSplat BACON Grid GT
2D Spheres
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
2D Bandlimited
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
2D Sierpinski
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
2D Star
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
3D Spheres
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
3D Bandlimited
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
DIV2K
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
3D Dragon Occ. Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
3D Dragon Surf. Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Figure 2: Qualitative overfitting results. Visualizations of each model on each overfitting task with $1\times 10^{4}$ parameters, roughly $1\%$ of the pixels/voxels in the original 2D and 3D signals. For 3D signals, a slice is visualized. For synthetic signals with a bandwidth parameter, bandwidth 0.5 is shown. GSplat is restricted to 2D signals. Full visualizations varying model size and signal bandwidth are provided in Section 5.6. Different parameterizations induce different characteristic qualitative compression artifacts.
Figure 3: Overfitting Synthetic Signals: INR and INR - Grid Heatmaps. Red indicates regimes where other models outperform the Grid baseline; blue indicates regimes dominated by the Grid baseline. Each first row shows absolute PSNR values, while each second row shows the PSNR gap relative to the Grid baseline (i.e., PSNR - Grid PSNR). See Section 3.1 for in-depth discussion.
Figure 4: Overfitting Capacity and Computational Efficiency. (a) Overfitting capacity of different models evaluated on the 2D DIV2K and 3D Stanford Dragon datasets. PSNR trends indicate how well each model fits the training data as a function of model size (top) as well as relative performance compared to the Grid baseline (bottom). (b) Computational efficiency analysis, showing inference and training times for each model. While WIRE achieves superior overfitting capacity in some cases, it requires significantly more computation time—approximately 10×10$\times$10 × that of the next slowest model.
2.2 Synthetic Signals
We generate diverse 2D and 3D synthetic signals to test how well different models can overfit to representative components found in natural signals: Spheres, Bandlimited signals, Sierpinski fractals, and a radial Star Target. Each synthetic signal has approximately $1\times 10^{6}$ total spatial values, usually arranged as resolution $1000^{2}$ for 2D signals and $100^{3}$ for 3D signals. This ensures that our model parameter budgets (from $1\times 10^{4}$ to $3\times 10^{6}$) focus on the compressive regime while also including some overparameterized models, and that the level of compression is independent of the signal dimension (2D or 3D). Each type of synthetic signal is designed with a variable scale parameter representing qualitative or quantitative bandwidth on a scale from $0.1$ to $0.9$, so that we can learn how well different models fit signals with different bandwidths. These signals are summarized in the top half of Tables 1, 1 and 2. More detail on our synthetic signals may be found in Section 5.3.
2.3 Real Signals
In addition to synthetic experiments, we evaluate each model on diverse tasks involving real-world 2D images and 3D volumes. These experiments assess the models’ ability to overfit, generalize, and solve inverse problems with natural signals. We use 10 images from DIV2K [27] to test image overfitting, 4×4$\times$4 × super-resolution for height and width, and denoising at two different levels of Gaussian noise. We use a 2D chest CT scan from Clark et al. [28] to evaluate performance in a classic underdetermined inverse problem, in which the task is to recover an image from undersampled X-ray projections. We evaluate volumetric performance using the occupancy function and surface of the 3D Stanford Dragon [29]. On both of these 3D signals we test both volume overfitting and super-resolution. Our real-data signals and tasks are summarized in the bottom half of Table 1, with further details provided in Section 5.4.
2.4 Evaluation Metrics
We use peak signal-to-noise ratio (PSNR) as the primary metric, with higher PSNR values indicating better reconstruction fidelity. All quantitative results evaluated using PSNR, SSIM, and LPIPS for 2D and PSNR and IOU for 3D can be found in the appendix. Note that our Grid baseline model is unregularized for overfitting tasks but uses total variation (TV) regularization for noisy or underdetermined tasks. Additional implementation details may be found in Section 5.5.
3 Results
Our experiments span overfitting tasks for synthetic and real signals, as well as more generalization-oriented tasks such as computed tomography (CT), denoising, and super-resolution (SR).
3.1 Overfitting
We first analyze how well each model can overfit 2D and 3D signals with a fixed parameter budget. Our goal is to quantify how representation capacity scales with model size: a successful overfitting indicates that the model can capture the signal’s complexity within the given parameter constraints.
Qualitative Evaluation.
Figure 2 visualizes the outputs of all models on all overfitting tasks, for the most compressive setting where each model is constrained to at most $1\times 10^{4}$ parameters, roughly $1\%$ of the raw signal sizes. For synthetic signals with variable “bandwidth” we show an intermediate-complexity signal (bandwidth=$0.5$). These visualizations illustrate the types of artifacts induced by compression with different signal representations. For example, Fourier Feature Networks [15] can induce aliasing-like artifacts. SIREN [14] and WIRE [16] work well for many signals but introduce texture artifacts for others, especially in 3D, while BACON suffers texture artifacts on all signals at this extreme compression level. Surprisingly, WIRE struggles to fit solid 3D Spheres but excels at representing a solid 3D dragon. GA-Planes [26] tends to yield the highest quality overfitting results at this most compressed scale, except for the highest-frequency portion of the Star Target and pure bandlimited signals, for which the grid excels. Under this extreme compression, Instant-NGP [22] suffers noisy artifacts on all signals, likely due to extensive hash collisions that are a byproduct of small hash tables. GSplat tends to represent a subset of each signal well, while failing to capture other portions, perhaps due to its random initialization and poorly-conditioned gradients for optimization. The simple interpolated Grid parameterization turns out to be quite a strong baseline across these diverse signals, with the expected limitation of blurring over super-Nyquist details that are beyond the grid resolution that is affordable under strong compression.
Comprehensive visualizations including all signal bandwidths and all model sizes are provided in Section 5.6. We specifically highlight Figure 10 which shows visual error maps for each model on the 2D Star Target. This allows us to visualize the distribution of modeling errors with respect to the scale of local image features, and how these errors change with model size. For example, errors in the Grid baseline are, as expected, concentrated at sharp edges; the Grid is the only model that successfully reconstructs the smooth background even at the smallest model size. We also observe that all models suffer the largest reconstruction errors towards the center of the Star Target, where the local signal frequency is highest. The size of this high-error region decays with increasing model size, demonstrating that INRs as well as hybrid, discrete, and grid representations all have an inherent bandwidth that grows with model size.
Quantitative Evaluation.
Figure 3 and Figure 4 report PSNR for each model on each overfitting task. Figure 3 focuses on synthetic signals with a notion of bandwidth, and stratifies performance within each heatmap according to model size ($y$ axis) and effective signal bandwidth ($x$ axis). For each signal, the top row of heatmaps reports the PSNR achieved by each model, while the second row reports the difference in PSNR between the INR, hybrid, or discrete method and the Grid baseline for each experiment. This second row allows rapid identification of the regimes (in red) in which INR, hybrid, or discrete methods can outperform the Grid baseline, versus regimes (in blue) where the Grid is the best-performing representation. First, we observe that several models, notably GA-Planes and Instant-NGP as well as some Fourier Feature Networks, can outperform the Grid baseline at intermediate model sizes on the synthetic signals with constant-value regions and sharp edges (Spheres, Sierpinski, and Star Target). We posit that these signals are easier to fit with an adaptive representation due to their underlying lower-dimensional structure, as each of these signals has its complexity concentrated at either a 2D surface embedded in 3D space or a 1D edge embedded in 2D space. However, we also observe two key takeaways evident in 2D and 3D Bandlimited signals: (1) most models exhibit an inherent bandwidth that increases with model size, apparently following a power law since both the model size and signal bandlimit are on a log scale for this signal class, and (2) no model can reliably outperform the simple Grid baseline for 2D or 3D Bandlimited signals, regardless of model size. This is a strong and perhaps surprising negative result suggesting that INRs are not an all-purpose solution but are instead best suited to representing specific types of signals.
Figure 4 (a) reports quantitative overfitting performance on real 2D and 3D signals including images from DIV2K [27] as well as the 3D Stanford Dragon [29] posed as occupancy and surface fitting tasks. While Grid outperforms other methods on 2D image overfitting, INRs exhibit competitive performance in 3D overfitting tasks—perhaps because, like our synthetic Spheres signal, the 3D Dragon consists of constant regions and sharp edges with underlying lower-dimensional structure. In particular, GA-Planes and WIRE demonstrate strong compression capabilities in the 3D object-fitting task, highlighting a potential use case for INR development and applications.
Figure 4 (b) compares inference and training times for different models on the 2D Star Target across various parameter budgets. The simple Grid model with bicubic interpolation provides the fastest inference and training times, pure INRs require the most computation time, and hybrid models (GA-Planes, Instant-NGP) and discrete (GSplat) models are in between. We note that the training and inference times for the Grid and hybrid models are in principle independent of model size, with some caveats depending on batch size and feature dimension (which explain the increasing trend for GA-Planes). For pure INRs there is an architecturally inherent increase in computation time with model size.
3.2 Generalization and Inverse Problems
FFN SIREN WIRE GA-Planes Instant-NGP GSplat BACON Grid GT
CT
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
$\epsilon=0.1$
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
SR
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Occ. SR
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Surf. SR
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Figure 5: Qualitative Generalization and Inverse Problems Results. In CT reconstruction, Grid with TV regularization achieves the best results. Experiments on the DIV2K dataset are zoomed in to highlight image details. For image denoising and super-resolution, WIRE produces sharp images but introduces texture artifacts. GA-Planes outperforms other methods in volume and surface super-resolution. Artifacts in each model’s output throughout the inverse problem tasks reveal their inherent structural biases (e.g., sinusoidal artifacts in FFNs and SIREN, line artifacts in GA-Planes).
                    CT               Denoising $\epsilon=0.1$         DIV2K Super Resolution   3D Dragon Occupancy Super Resolution     3D Dragon Surface Super Resolution
PSNR Gap (dB)          PSNR (dB)
Figure 6: Quantitative Generalization and Inverse Problems Results. PSNR vs. model size trends indicate how well each model performs each generalization task as a function of model size (top) as well as relative performance compared to the TV-regularized Grid baseline (bottom). WIRE and GA-Planes perform best at highly compressed 3D super-resolution, but the simple Grid baseline generally outperforms other signal representation methods across these tasks.
Visual and quantitative results for computed tomography, denoising, and super-resolution are presented in Figure 5 and Figure 6, respectively. For all of these tasks involving natural 2D signals, we find that the simple Grid baseline with total variation (TV) regularization is optimal across all model sizes. However, on our 3D Dragon occupancy and surface super-resolution tasks, GA-Planes, WIRE, and to some extent SIREN outperform the Grid at the smallest model sizes. As in our overfitting experiments, this suggests that some INRs and hybrid models may be best suited to representing and compressing signals with constant regions and sharp edges characteristic of lower-dimensional structure. Comprehensive visualizations and quantitative metrics including an additional denoising task with $\epsilon=0.05$ are provided in Section 5.7.
4 Discussion
Our results shed light on the implicit biases of different signal representation methods: how each method allocates its limited capacity across signals with varying characteristics, how this capacity scales with model size, and what visual artifacts are induced by compression with each method. Our experiments span 2D and 3D synthetic and real datasets as well as diverse tasks including overfitting, tomography, denoising, and super-resolution. By stratifying performance across signal bandwidth, model size, and signal type, we provide insights into the strengths and limitations of implicit neural representations (INRs), hybrid models, discrete models, and a simple yet surprisingly strong grid-based baseline.
Our experiments reveal several key findings. First, for bandlimited signals, we observe a power-law relationship between model size and bandwidth across nearly all models, providing a tool to quantify the inherent “resolution” of INRs as a function of parameter count. For 2D and 3D bandlimited signals, a simple interpolated grid representation consistently outperforms other representations in both computational efficiency and reconstruction quality, using equal parameter budgets. In most inverse problems, interpolated grids with regularization again offer superior performance, training faster and achieving higher quality results than other models with the same number of parameters. However, INRs exhibit advantages in specific scenarios for both overfitting and generalization tasks in 2D and 3D, specifically when the target signal has an underlying lower-dimensional strucutre, such as constant-value regions with sharp edges.
Overall, our results emphasize that a simple grid with interpolation remains the most practical and effective choice for a wide range of applications, offering simplicity, interpretability, computational efficiency, and in many cases optimal reconstruction quality. However, INRs provide distinct advantages in representing specific types of signals, specifically those with underlying lower-dimensional structure such as object edges or surfaces. By analyzing these trade-offs, our study offers practical guidelines for selecting the most suitable representation method for different signal types and applications and for guiding the development of future signal representations.
Acknowledgements
This work was supported in part by the NSF Mathematical Sciences Postdoctoral Research Fellowship under award number 2303178. Any opinions, findings, and conclusions or recommendations expressed in this material are those of the authors and do not necessarily reflect the views of the National Science Foundation.